Make a quadratic equation whose roots are -2 + sqrt( 5 ) and -2 - sqrt( 5 ).
Answer by edjones(8007)

x=-2+sqrt(2), x=-2-sqrt(2)
(x+2)^2=2
x^2+4x+4-2=0
x^2+4x+2=0
.
Ed